OpenStudy (anonymous):

How would you solve this equation: 64^x=4^x+4 (the +4 is with the x, the exponent over the 4 is x+4)

OpenStudy (anonymous):

\[64^{x} = 4^{x+4}\] <--- like that?

OpenStudy (anonymous):

Yes, like that.

OpenStudy (anonymous):

okay so thing of how u can simplify 64 how many times does 4 go into 16?

OpenStudy (anonymous):

* 64

OpenStudy (anonymous):

It goes into it 16 times. What I have so far is \[4^{3+x}=4^{x+4}\]

OpenStudy (anonymous):

did u get it? Do you want me to walk u through it or do you want me to just post the steps

OpenStudy (anonymous):

Walk me through it would be the best way.

OpenStudy (anonymous):

PERFECT since the bases are the same, you cancel them out and it becomes 3+x = x+4

OpenStudy (anonymous):

Yeah, I got that.

OpenStudy (anonymous):

Now it's simple algebra.. do you know what to do now, or do you want me to continue?

OpenStudy (anonymous):

Continue please, so that I don't mess up.

OpenStudy (anonymous):

so now you have to get the xes to 1 side and the numbers to the otherside so... 4-3= 1 x-x =0 0, thus there is No solution..... to check : http://www.wolframalpha.com/input/?i=64%5Ex%3D4%5Ex%2B4

OpenStudy (anonymous):

(4^3)^x=4^(x+4) 4^(3x)=4^(x+4) 3x=x+4 2x=4 x=2

OpenStudy (anonymous):

You added +4 to 4^x, it's actaully \[4^{x+4}\] not \[4^{x}+4\]

OpenStudy (anonymous):

ohhhh wait hold up

OpenStudy (anonymous):

\[4^{3x}=4^{x+4}\]

OpenStudy (anonymous):

that's where ur wrong you did 3+x in stead of 3x

OpenStudy (anonymous):

When I tried it, I got x=1. These are the steps I did: \[4^{3}=4^{x+4}\] \[3^{x}=x+4\] \[2x=4\] x=2

OpenStudy (anonymous):

so the real answer is ... 3x=x+4 3x-x= 2x 2x=4 divide by 2 x=2

OpenStudy (anonymous):

Yes, we were both wrong at certain points. Thanks.

OpenStudy (anonymous):

lol ur welcome